# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 6.1 Path decomposition
We will here interpret the coordinates of the matrix $A_{\ell,\#}$ as paths in a graph.
@@
By convention we assume that the sum over an empty set is 00 and the product over an empty set is $1$. With the previous definition we have
$(A_{\ell})_{xy}=\sum_{\gamma\in\Gamma_{xy}^{\ell}}\prod_{t=1}^{T_{\gamma}}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$
@@
  • We call a coincidence a sequence $c=(x_{1},y_{1},...,x_{T},y_{T},x_{T+1})$ such that $(x_{2},...,x_{T+1})$ are pairwise distinct and such that there exists $(\ell_{1},...,\ell_{T})$ a $T$-tuple of integers verifying $\sum_{t}\ell_{t}+T-1\leq\ell$, $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,y_{T},\ell_{T},x_{T+1})$ is a path and $(B_{1}^{\ell_{1}}B_{2}^{\ell_{2}}\cdots B^{\ell_{T}}_{T})_{x_{1},x_{T+1}}>0$ where $B_{1},...,B_{T}\in\{Q,Q^{\ast}\}$ are alternate terms (i.e. $B_{t+1}=B_{t}^{\ast}$).
  • We say that a path $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,x_{T},y_{T},\ell_{T},x_{T+1})$ is tangle-free if the sequence $(x_{1},y_{1},x_{2},...,x_{T},y_{T},x_{T+1})$ contains, as a subsequence, at most one coincidence. We denote by $\mathrm{F}^{k}$ (resp. $\mathrm{F}_{xy}^{k}$) the subset of $\Gamma^{k}$ (resp. $\Gamma_{xy}^{k}$) of path of length $k$ (resp. path of length $k$ from $x$ to $y$) that are tangle-free.
  • The pair $(M,Q)$ is $\ell$-tangle-free if for all $k\in[\ell]$, for all $\gamma=(x_{1},y_{1},\ell_{1},...,y_{T},\ell_{T},x_{T+1})\in\Gamma^{k}\backslash\mathrm{F}^{k}$ we have
@@
$A_{\ell}=A^{(\ell)}$
where we defined $A^{(\ell)}$ by
$(A^{(\ell)})_{xy}:=\sum_{\gamma\in\mathrm{F}^{\ell}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}},$ (18)
for $x,y\in[N]$. Similarly we define the matrix $\underline{A}^{(\ell)}$ by its coordinates
$(\underline{A}^{(\ell)})_{xy}:=\sum_{\gamma\in\mathrm{F}^{\ell}_{xy}}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}$ (19)
@@
$(A^{(k)})_{xy}$ $=\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}=\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}(A^{(\vec{\ell})})_{xy}$ (21)
@@
$(A^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$ (22)
@@
$(\underline{A}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$ (23)
@@
$(A^{(\vec{\ell})})_{xy}=(\underline{A}^{(\vec{\ell})})_{xy}+\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\sum_{k=1}^{T}\left(\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right)\frac{(Q^{\ell_{k}})_{y_{k}x_{k+1}}}{N}\prod_{t=k+1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$
@@
$(A^{(\vec{\ell})})_{xy}=(\underline{A}^{(\vec{\ell})})_{xy}+\sum_{k=1}^{T}\frac{1}{N}\underbrace{\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}}_{(S^{(\vec{\ell})})_{xy}}.$
@@
$\mathrm{F}_{xy}^{\vec{\ell}}\sqcup\mathrm{T}_{xy}^{k,\vec{\ell}}=\bigsqcup_{x_{k}\in[N]}\bigsqcup_{y_{k}\in[N]}\bigsqcup_{x_{k+1}\in[N]}\{\gamma=(\gamma^{\prime},y_{k},\ell_{k},\gamma^{\prime\prime})\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \gamma^{\prime}\in\mathrm{F}^{\vec{\ell}(k)}_{xx_{k}},\leavevmode\nobreak\ \gamma^{\prime\prime}\in\mathrm{F}^{\vec{\ell}(k,T)}_{x_{k+1}y}\}$
@@
$(R_{k}^{(\vec{\ell})})_{xy}:=\sum_{\gamma\in\mathrm{T}^{k,\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}.$ (24)
@@
$(S^{(\vec{\ell})})_{xy}$ $=\sum_{x_{k},y_{k},x_{k+1}=1}^{N}\sum_{\begin{subarray}{c}\gamma^{\prime}\in\mathrm{F}_{xx_{k}}^{\vec{\ell}(k)}\\ \gamma^{\prime\prime}\in\mathrm{F}^{\vec{\ell}(k,T)}_{x_{k+1}y}\end{subarray}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}-(R_{k}^{(\vec{\ell})})_{xy}$
$=\sum_{y_{k}=1}^{N}\sum_{x_{k}=1}^{N}\left(\underline{P}^{(\vec{\ell}(k))}\right)_{xx_{k}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\sum_{x_{k+1}=1}^{n}\left(A^{(\vec{\ell}(k,T))}\right)_{x_{k+1}y}-(R_{k}^{(\vec{\ell})})_{xy}$
$=\left(A^{(\vec{\ell}(k))}Q^{\ell_{k}}(1\otimes 1)A^{(\vec{\ell}(k,T))}\right)_{xy}-(R_{k}^{(\vec{\ell})})_{xy}$
@@
$A^{(\vec{\ell})}=\underline{A}^{(\vec{\ell})}+\frac{1}{N}\sum_{k=1}^{T-1}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}(1\otimes 1)A^{(\vec{\ell}(k,T))}-\frac{1}{N}\sum_{k=1}^{T}R_{k}^{(\vec{\ell})}.$
@@
$A^{(\ell)}=\underline{A}^{(\ell)}+\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\left[\sum_{k=1}^{T}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\vec{\ell}(k,T))}-\frac{1}{N}\sum_{k=1}^{T}R_{k}^{(\vec{\ell})}\right].$
To estimate the operator norm of $A^{(\ell)}|_{\boldsymbol{1}^{\perp}}$ we first show that the terms with $\boldsymbol{1}\otimes\boldsymbol{1}$ in the previous formula have no contribution. For $1\leq T\leq\ell$ and $\vec{\ell}\in\mathcal{L}(T,\ell)$, we set the notations $\ell(k):=\sum_{t=1}^{k-1}\ell_{t}+(k-1)-1$ and $\ell(k,T):=\sum_{t=k+1}^{T}\ell_{t}+T-k-1$. We then have $\ell(k)+\ell(k,T)+\ell_{k}=\ell-2$ and we can partition $\mathcal{L}(\ell,T)$ as follows
$\mathcal{L}(\ell,T)=\bigsqcup_{\ell_{k}=0}^{\ell-T+1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bigsqcup_{\ell(k)=0}^{\ell-\ell_{k}-2}\mathcal{L}(\ell(k),k-1)\times\{\ell_{k}\}\times\mathcal{L}(\ell-\ell_{k}-2-\ell(k),T-k)$
@@
$\phi(\ell)$ $:=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\vec{\ell}(k,T))}$
$=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\ell_{k}=0}^{\ell}\sum_{\ell(k)=0}^{\ell-\ell_{k}-2}\sum_{\begin{subarray}{c}\vec{\ell}(k)\in\mathcal{L}(\ell(k),k-1)\\ \vec{\ell}(k,T)\in\mathcal{L}(\ell-\ell_{k}-\ell(k)-2,T-k)\end{subarray}}\underline{A}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\vec{\ell}(k,T))}$
$=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\ell_{k}=0}^{\ell}\sum_{\ell(k)=0}^{\ell-\ell_{k}-2}\underline{A}^{(\ell(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A^{(\ell-\ell_{k}-\ell(k)-2)}.$
[…]
$\boldsymbol{1}\otimes\boldsymbol{1}A^{(\ell^{\prime})}w=\boldsymbol{1}\otimes\boldsymbol{1}A_{\ell^{\prime}}w=0.$
@@
$A^{(\ell)}=\underline{A}^{(\ell)}+\phi(\ell)-\frac{1}{N}\sum_{k=1}^{\ell}R^{(\ell)}_{k}.$
To have consistent writings between the operators $R_{k}^{(\ell)}$ and $A^{(\ell)}$, for $1\leq k\leq\ell$ and $x,y\in[N]$ we set
@@
$(R_{k}^{(\ell)})_{xy}=\sum_{\gamma\in\mathrm{T}_{xy}^{k,\ell}}\left(\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right)(Q^{\ell_{k}})_{y_{k}x_{k+1}}\left(\prod_{t=k+1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right).$
When it comes to the case #$\#=r$, all the previous reasoning holds if now one considers $A_{r}=(1-r)M+rQ$ and replace respectively (22) and (23) with
$(A_{r}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}(1-r)^{T}\prod_{t=1}^{T}M_{x_{t}y_{t}}(\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ (\underline{A}_{r}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}(1-r)^{T}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}.$
@@
$(R_{k,r}^{(\vec{\ell})})_{xy}:=(1-r)^{T-1}\sum_{\gamma\in\mathrm{T}^{k,\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}\{rQ\}^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}\{rQ\}^{\ell_{v}})_{y_{v}x_{v+1}}.$ (25)
@@
[…]
$\|A_{\#}^{\ell}|_{\boldsymbol{1}^{\perp}}\|\leq\|\underline{A}_{\#}^{(\ell)}\|+\frac{1}{N}\sum_{k=1}^{\ell}\|R^{(\ell)}_{k,\#}\|.$

### v4

@@ 7.2 Path decomposition
We will here interpret the coordinates of the matrix $A_{\ell,\mathrm{x}}$ as paths in a graph.
@@
By convention we assume that the sum over an empty set is 0 and the product over an empty set is $1$. With the previous definition we have
$(A_{\ell,\mathrm{s}})_{xy}=\sum_{\gamma\in\Gamma_{xy}^{\ell}}\prod_{t=1}^{T_{\gamma}}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$
@@
  • We call a coincidence a sequence $c=(x_{1},y_{1},...,x_{T},y_{T},x_{T+1})$ such that $(x_{2},...,x_{T+1})$ are pairwise distinct and such that there exists $(\ell_{1},...,\ell_{T})$ a $T$-tuple of integers verifying $\sum_{t}\ell_{t}\leq\ell$, $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,y_{T},\ell_{T},x_{T+1})$ is a path and $(B_{1}^{\ell_{1}}B_{2}^{\ell_{2}}\cdots B^{\ell_{T}}_{T})_{x_{1},x_{T+1}}>0$ where $B_{1},...,B_{T}\in\{Q,Q^{\ast}\}$ are alternate terms (i.e. $B_{t+1}=B_{t}^{\ast}$).
  • We say that a path $\gamma=(x_{1},y_{1},\ell_{1},x_{2},...,x_{T},y_{T},\ell_{T},x_{T+1})$ of length $\ell$ is tangle-free if the sequence $(x_{1},y_{1},x_{2},...,x_{T},y_{T},x_{T+1})$ contains, as a subsequence, at most one coincidence. We denote by $\mathrm{F}^{k}$ (resp. $\mathrm{F}_{xy}^{k}$) the subset of $\Gamma^{k}$ (resp. $\Gamma_{xy}^{k}$) of path of length $k$ (resp. path of length $k$ from $x$ to $y$) that are tangle-free.
[…]
@@
$A_{\ell,\mathrm{s}}=A_{\mathrm{s}}^{(\ell)}$
where we defined $A_{\mathrm{s}}^{(\ell)}$ by
$(A_{\mathrm{s}}^{(\ell)})_{xy}:=\sum_{\gamma\in\mathrm{F}^{\ell}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}},$ (21)
for $x,y\in[N]$. Similarly we define the matrix $\underline{A}_{\mathrm{s}}^{(\ell)}$ by its coordinates
$(\underline{A}_{\mathrm{s}}^{(\ell)})_{xy}:=\sum_{\gamma\in\mathrm{F}^{\ell}_{xy}}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}$ (22)
@@
$(A_{\mathrm{s}}^{(k)})_{xy}$ $=\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}=\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}(A^{(\vec{\ell})})_{xy}$ (24)
@@
$(A_{\mathrm{s}}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$ (25)
@@
$(\underline{A}_{\mathrm{s}}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$ (26)
@@
$(A_{\mathrm{s}}^{(\vec{\ell})})_{xy}=(\underline{A}_{\mathrm{s}}^{(\vec{\ell})})_{xy}+\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\sum_{k=1}^{T}\left(\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right)\frac{(Q^{\ell_{k}})_{y_{k}x_{k+1}}}{N}\prod_{t=k+1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}.$
@@
$(A_{\mathrm{s}}^{(\vec{\ell})})_{xy}=(\underline{A}_{\mathrm{s}}^{(\vec{\ell})})_{xy}+\sum_{k=1}^{T}\frac{1}{N}\underbrace{\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}}_{(S^{(\vec{\ell})})_{xy}}.$
@@
$\mathrm{F}_{xy}^{\vec{\ell}}\sqcup\mathrm{T}_{xy}^{k,\vec{\ell}}=\bigsqcup_{x_{k}\in[N]}\bigsqcup_{y_{k}\in[N]}\bigsqcup_{x_{k+1}\in[N]}\{\gamma=(\gamma^{\prime},y_{k},\ell_{k},\gamma^{\prime\prime})\penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \gamma^{\prime}\in\mathrm{F}^{\vec{\ell}(k)}_{xx_{k}},\penalty 10000\ \gamma^{\prime\prime}\in\mathrm{F}^{\vec{\ell}(k,T)}_{x_{k+1}y}\}$
@@
$(R_{k,\mathrm{s}}^{(\vec{\ell})})_{xy}:=\sum_{\gamma\in\mathrm{T}^{k,\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}.$ (27)
@@
$(S^{(\vec{\ell})})_{xy}$ $=\sum_{x_{k},y_{k},x_{k+1}=1}^{N}\sum_{\begin{subarray}{c}\gamma^{\prime}\in\mathrm{F}_{xx_{k}}^{\vec{\ell}(k)}\\ \gamma^{\prime\prime}\in\mathrm{F}^{\vec{\ell}(k,T)}_{x_{k+1}y}\end{subarray}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}(Q^{\ell_{v}})_{y_{v}x_{v+1}}-(R_{k,\mathrm{s}}^{(\vec{\ell})})_{xy}$
$=\sum_{y_{k}=1}^{N}\sum_{x_{k}=1}^{N}\left(\underline{P}^{(\vec{\ell}(k))}\right)_{xx_{k}}(Q^{\ell_{k}})_{y_{k}x_{k+1}}\sum_{x_{k+1}=1}^{n}\left(A_{\mathrm{s}}^{(\vec{\ell}(k,T))}\right)_{x_{k+1}y}-(R_{k,\mathrm{s}}^{(\vec{\ell})})_{xy}$
$=\left(A_{\mathrm{s}}^{(\vec{\ell}(k))}Q^{\ell_{k}}(1\otimes 1)A_{\mathrm{s}}^{(\vec{\ell}(k,T))}\right)_{xy}-(R_{k,\mathrm{s}}^{(\vec{\ell})})_{xy}$
@@
$A_{\mathrm{s}}^{(\vec{\ell})}=\underline{A}_{\mathrm{s}}^{(\vec{\ell})}+\frac{1}{N}\sum_{k=1}^{T-1}\underline{A}_{\mathrm{s}}^{(\vec{\ell}(k))}Q^{\ell_{k}}(1\otimes 1)A_{\mathrm{s}}^{(\vec{\ell}(k,T))}-\frac{1}{N}\sum_{k=1}^{T}R_{k,\mathrm{s}}^{(\vec{\ell})}.$
@@
$A_{\mathrm{s}}^{(\ell)}=\underline{A}_{\mathrm{s}}^{(\ell)}+\sum_{T=1}^{\ell}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\left[\sum_{k=1}^{T}\underline{A}_{\mathrm{s}}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A_{\mathrm{s}}^{(\vec{\ell}(k,T))}-\frac{1}{N}\sum_{k=1}^{T}R_{k,\mathrm{s}}^{(\vec{\ell})}\right].$
To estimate the operator norm of $A_{\mathrm{s}}^{(\ell)}|_{\boldsymbol{1}^{\perp}}$ we first show that the terms with $\boldsymbol{1}\otimes\boldsymbol{1}$ in the previous formula have no contribution. For $1\leq T\leq\ell$ and $\vec{\ell}\in\mathcal{L}(T,\ell)$, we set the notations $\ell(k):=\sum_{t=1}^{k-1}\ell_{t}+(k-1)-1$ and $\ell(k,T):=\sum_{t=k+1}^{T}\ell_{t}+T-k-1$. We then have $\ell(k)+\ell(k,T)+\ell_{k}=\ell-2$ and we can partition $\mathcal{L}(\ell,T)$ as follows
$\mathcal{L}(\ell,T)=\bigsqcup_{\ell_{k}=0}^{\ell-T+1}\penalty 10000\ \penalty 10000\ \penalty 10000\ \bigsqcup_{\ell(k)=0}^{\ell-\ell_{k}-2}\mathcal{L}(\ell(k),k-1)\times\{\ell_{k}\}\times\mathcal{L}(\ell-\ell_{k}-2-\ell(k),T-k)$
@@
$\phi(\ell)$ $:=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\vec{\ell}\in\mathcal{L}(T,\ell)}\underline{A}_{\mathrm{s}}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A_{\mathrm{s}}^{(\vec{\ell}(k,T))}$
$=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\ell_{k}=0}^{\ell}\sum_{\ell(k)=0}^{\ell-\ell_{k}-2}\sum_{\begin{subarray}{c}\vec{\ell}(k)\in\mathcal{L}(\ell(k),k-1)\\ \vec{\ell}(k,T)\in\mathcal{L}(\ell-\ell_{k}-\ell(k)-2,T-k)\end{subarray}}\underline{A}_{\mathrm{s}}^{(\vec{\ell}(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A_{\mathrm{s}}^{(\vec{\ell}(k,T))}$
$=\sum_{T=1}^{\ell}\sum_{k=1}^{T}\sum_{\ell_{k}=0}^{\ell}\sum_{\ell(k)=0}^{\ell-\ell_{k}-2}\underline{A}_{\mathrm{s}}^{(\ell(k))}Q^{\ell_{k}}\boldsymbol{1}\otimes\boldsymbol{1}A_{\mathrm{s}}^{(\ell-\ell_{k}-\ell(k)-2)}.$
On the event $(M,Q)$ $\ell$-tangle free, for all $0\leq k\leq\ell$ and for all $w\in\boldsymbol{1}^{\perp}$ we have
$\boldsymbol{1}\otimes\boldsymbol{1}A_{\mathrm{s}}^{(k)}w=\boldsymbol{1}\otimes\boldsymbol{1}A_{k,\mathrm{s}}w=0.$
@@
$A_{\mathrm{s}}^{(\ell)}=\underline{A}_{\mathrm{s}}^{(\ell)}+\phi(\ell)-\frac{1}{N}\sum_{k=1}^{\ell}R^{(\ell)}_{k,\mathrm{s}}$.
To have consistent writings between the operators $R_{k,\mathrm{s}}^{(\ell)}$ and $A_{\mathrm{s}}^{(\ell)}$, for $1\leq k\leq\ell$ and $x,y\in[N]$ we set
@@
$(R_{k,\mathrm{s}}^{(\ell)})_{xy}=\sum_{\gamma\in\mathrm{T}_{xy}^{k,\ell}}\left(\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right)(Q^{\ell_{k}})_{y_{k}x_{k+1}}\left(\prod_{t=k+1}^{T}M_{x_{t}y_{t}}(Q^{\ell_{t}})_{y_{t}x_{t+1}}\right).$
When it comes to the case $\mathrm{x}=\mathrm{c}$, all the previous reasoning holds if now one considers $A_{\mathrm{c}}=(1-r)M+rQ$ and replace respectively (25) and (26) with
$(A_{\mathrm{c}}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}(1-r)^{T}\prod_{t=1}^{T}M_{x_{t}y_{t}}(\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ (\underline{A}_{\mathrm{c}}^{(\vec{\ell})})_{xy}=\sum_{\gamma\in\mathrm{F}^{\vec{\ell}}_{xy}}(1-r)^{T}\prod_{t=1}^{T}\underline{M}_{x_{t}y_{t}}(\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}.$
@@
$(R_{k,\mathrm{c}}^{(\vec{\ell})})_{xy}:=(1-r)^{T-1}\sum_{\gamma\in\mathrm{T}^{k,\vec{\ell}}_{xy}}\prod_{t=1}^{k-1}\underline{M}_{x_{t}y_{t}}\{rQ\}^{\ell_{t}})_{y_{t}x_{t+1}}\{rQ\}^{\ell_{k}})_{y_{k}x_{k+1}}\prod_{v=k+1}^{T}M_{x_{v}y_{v}}\{rQ\}^{\ell_{v}})_{y_{v}x_{v+1}}.$ (28)
@@
On the event $(M,Q)$ $\ell$-tangle free we have
$\|A_{\mathrm{x}}^{\ell}|_{\boldsymbol{1}^{\perp}}\|\leq\|\underline{A}_{\mathrm{x}}^{(\ell)}\|+\frac{1}{N}\sum_{k=1}^{\ell}\|R^{(\ell)}_{k,\mathrm{x}}\|$.